OpenStudy (anonymous):

Find f '(−3), if f(x) = (2x2 − 7x)(−x2 − 7). Round your answer to the nearest integer. Use the hyphen symbol, -, for negative values.

OpenStudy (mathmale):

this function is presented as a product. You could either multiply out the two given terms and then differentiate the results, OR you could apply the Product Rule, the Power Rule and several others.

OpenStudy (mathmale):

Anyone care to post and explain the Product Rule for differentiation?

OpenStudy (anonymous):

Give me a second

OpenStudy (anonymous):

\[(f*g)\prime (x) = f(x)\prime* g(x) + f(x)* g(x)\prime \] is the product rule

OpenStudy (mathmale):

if f(x) = (2x2 − 7x)(−x2 − 7, which is in the form of a product, the product of (2x^2 − 7x) and (−x2 − 7), use the product rule to find the derivative of f(x). Hint: let u= (2x^2 − 7x) and v= (−x^2 − 7). Next, find u'(x) and v'(x). Last, substitute your u, v, u' and v' into the Product Rule. @csheveland has typed this in for you (see above).